8.3 4) the claim is that for 12 am body temperature the mean is >98.6 degrees...

8.3
4) the claim is that for 12 am body temperature the mean is >98.6 degrees Fahrenheit. the sample size is n= 7 and the test statistic is t= 1.649. round to 3 decimal places as needed.
p value =_______
Answer #1

Degrees of freedom = n - 1 = 7 - 1 = 6

Hence,

p - Value = P(t6 > 1.649) = 0.075